A 50 kg skater B collides with a 40 kg skater A initially at rest on ice. They move together after the collision at 5.6 m/s find the velocity of the skater B before the collision


1
Expert's answer
2021-04-13T06:35:57-0400
"m_Av_A+m_Bv_B=(m_A+m_B)v_c,""0+m_Bv_B=(m_A+m_B)v_c""v_B=\dfrac{(m_A+m_B)v_c}{m_B},""v_B=\dfrac{(40\ kg+50\ kg)\cdot5.6\ \dfrac{m}{s}}{50\ kg}=10.1\ \dfrac{m}{s}."